OpenStudy (anonymous):

3x^2-7x-5=0 solutions? and x-intercepts?

OpenStudy (lgbasallote):

solution -> solve for x x-int _> just same

sam (.sam.):

use quadratic formula

OpenStudy (anonymous):

You kinda have to use the quadratic formula. \[-b ±\sqrt{b ^{2} - 4ac}\over2a\]

OpenStudy (anonymous):

I know I have to use the quadratic formula........the problem is using the formula

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

I am not sure how?

OpenStudy (anonymous):

a = 3 b =-7 and c = -5 You will get a complex answer.

OpenStudy (anonymous):

a = 3 b = -7 c = -5 Just plug in numbers. The form for a quadratic is ax^2 + bx + c That's where it comes in.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Im not sure how to calc x

OpenStudy (anonymous):

Your x will equal the complex number you get when you solve the quadratic equation..that is the x you are solving for.... x = quadratic equation.

OpenStudy (anonymous):

\[1/6(7-\sqrt{109},1/6(7+\sqrt{109}\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the x-intrecepts how do I write those?

OpenStudy (anonymous):

a = 3 b = -7 c = -5 (7 ±√109)/6 You can approximate it if you want, but these are the solutions. Last line.

OpenStudy (anonymous):

ordered pair for x-intr (-1,2)?